OpenStudy (anonymous):

Solve. logx - 1 16 = 4?‎

OpenStudy (anonymous):

\[\LARGE \log x-116=4\] \[\LARGE \log x=4+166\] \[\LARGE \log x=120\] is it like: \[\LARGE \log x-116=4\] or like: \[\LARGE \log (x-116)=4\] ?

OpenStudy (anonymous):

sorry it's 116 on the second line...

OpenStudy (anonymous):

x = 5 x = 4 x = 3 x = 2 ^ answer options. so im guessing the second way you did it is correct.

OpenStudy (anonymous):

\[\LARGE \log(x-116)=4\] has base 10 \[\LARGE \log_{10}(x-116)=4\] \[\LARGE x-116=10^4\] \[\LARGE x=10^4+116\] I would suggest you to recheck the given. Write them precisely... O_O (Or maybe I'm dumb, and I'm doing a mistake somewhere !)

OpenStudy (anonymous):

ohhh. i understand how to do it now, i believe you're doing it right i was just confused.

OpenStudy (anonymous):

is it 116 or a different number I see you've put a space between 1 16 ! .. or maybe is \[\LARGE \log_x(-116)=4\] but it doesn't make sense either... @kerrigan Tell me , I want to know why I can't do it :(

OpenStudy (anonymous):

there is a space between the 1 and the 16

OpenStudy (anonymous):

and that means? ... :O

OpenStudy (anonymous):

honestly, im a blonde. math isnt my thing. but how it is written is like this Solve. logx - 1 16 = 4?‎ the -1 in the problem is pushed down like under the rest of the equation... why? i have no idea.

OpenStudy (anonymous):

can you post a photo ? :A

OpenStudy (anonymous):

or draw it...

OpenStudy (anonymous):

yeah hold on.

OpenStudy (anonymous):